OpenStudy (bacon):

Please help Find the value of x so that the line containing (4, 8) and (2, -1) is perpendicular to the line containing (x, 2) and (-4, 5). Graph the lines.

OpenStudy (bacon):

How do i solve this step by step?

OpenStudy (anonymous):

sorry but i don't know how to solve that step by step

OpenStudy (anonymous):

first you get the equation of the line containing (4, 8) and (2,-1)...

OpenStudy (anonymous):

you are given two points already so you can plot the line....

OpenStudy (bacon):

ok, so the slope for that line would be 9/2

OpenStudy (bacon):

and then the slope for the second line would be -2/9 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yes of course

OpenStudy (bacon):

so then what would i do to find x?

OpenStudy (anonymous):

what would be the equation of the first line?

OpenStudy (anonymous):

you can determine it using point-slope formula...

OpenStudy (bacon):

so y-8 = 9/2(x-4)

OpenStudy (bacon):

would that be the equation for the first line?

OpenStudy (bacon):

actually would the end part be (x-(-4)

OpenStudy (anonymous):

yup better to write it in standard form.... Ax + By + C = 0

OpenStudy (bacon):

how would i do that?

OpenStudy (anonymous):

simplify your equation...

OpenStudy (anonymous):

A, B, and C are constants...

OpenStudy (anonymous):

y-8 = 9/2 (x-4) 2y-16=9x-36 9x-2y-36+16=0 9x-2y-20=0

OpenStudy (anonymous):

for the second line... it would be Bx + Ay + D=0

OpenStudy (anonymous):

in line1: A=9, B=-2, C=-20 for line2: A=-2, B=9, D still unknown...

OpenStudy (bacon):

hmmmm I am a bit lost, I'll just go and ask my teacher tomorrow

OpenStudy (bacon):

thanky ou fr trying though!

OpenStudy (anonymous):

line2 is -2x+9y+D=0

OpenStudy (anonymous):

you don't want to continue?

OpenStudy (anonymous):

we're almost there....

OpenStudy (bacon):

no, it is ok! Sorry for wasting your time :(

OpenStudy (anonymous):

the coefficients A and B for perpendicular lines are interchangeable...

OpenStudy (anonymous):

that is a rule...

OpenStudy (anonymous):

ok if that's what you want... :)